OpenStudy (anonymous):

I need to help finding the 50th and 100th term of the infinite series ½ + ⅔ + ¾ +⅘ +... i can't seem to find a common ratio among the terms...

OpenStudy (campbell_st):

thats because its not geometric the series appears to be \[\sum_{n=1}^{\infty} \frac{n}{n + 1}\] so substitute n = 50 to get the 50th term then n = 100 to get the 100th term.... hope it helps

OpenStudy (anonymous):

oh yeah, thanks. No wonder it wasnt making sense. thank you so much